OpenStudy (anonymous):

12-n, 12, 12+n What Is the average (arithmetic mean) of the 3 quantities in the list above? PLEASE HELP

OpenStudy (anonymous):

Av = 12 (12-n+12+12+n)/3 = 36/3 =12

OpenStudy (amistre64):

uh ... yeah. it tends to be the number in the middle

OpenStudy (amistre64):

if you have n to the left of you and n to the right of you; id pretty much say its a safe bet to say that you are in the middle